the average age of a student in graduate is normally distributed with a mean of 28 years of age and a standard deviation of 3 years of age. what percent of graduate students are at least 34 years old? possible answers are 1. 0,15% 2. 2.35% 3. 2.50% 4. 97.5% 5. None of the above
Answer #1

Given,

​​​​​​ = 28 , = 3

We convert this to standard normal as

P( X < x) = P( Z < x - / )

So,

P( X >= 34) = P( Z >= 34 - 28 / 3)

= P( Z >= 2)

= 1 - P( Z < 2)

= 1 - 0.9772

= 0.0228

= 2.28%

Answer: None of the above.